OpenStudy (anonymous):

4n-12=-24 What is n?

OpenStudy (anonymous):

n=-3

OpenStudy (jhannybean):

add 12 to both sides.

OpenStudy (anonymous):

Since -12 does not contain the variable variable A letter used to represent a number value in an expression or an equation. to solve for, move it to the right-hand side of the equation equation A mathematical statement that says two expressions have the same value; any number sentence with an = . by adding 12 to both sides. 4n=12-24 Subtract 24 from 12 to get -12. 4n=-12 Divide divide To perform the operation of division. each term term Any expression written as a product or quotient. Example: 2xy , , or in the equation equation A mathematical statement that says two expressions have the same value; any number sentence with an = . by 4. Simplify the left-hand side of the equation equation A mathematical statement that says two expressions have the same value; any number sentence with an = . by canceling the common factors factor One of two or more expressions that are multiplied together to get a product. .

OpenStudy (jhannybean):

\[\large 4n -12 +12 =-24 +12\] simplify. \[\large 4n = -12\]Now divide by 4 to both sides. \[\large \frac{4n}{4}= \frac{-12}{4}\] simplify. \[\large n=-3\]